An automatic gun fires ' n ' bullets with a speed of 1800 m/ sec in 3 seconds. The mass of each bullet is 25 gm. If 450N force is required to hold the gun in position.Find the number of bullets fired in each second.
1
Expert's answer
2017-04-03T11:00:07-0400
Force exerted by n bullets is 450 N (as the system remains stable).
Force exerted by n bullets: F=ma 450=25*10-3*a a=18000 m/s2
v=u+at
Initial velocity of a bullet is u=0 m/s
Final velocity of a bullet v=1800 m/s
1800=18000·t
t = 0.1 s (time taken for 1 bullet to be fired)
Number of bullets: 3/0.1=30.